# Boosting Data Analytics with Synthetic Volume Expansion

### 07 Jan 2024 v2

@@ 2.4 Generative Model and Knowledge Transfer
The Syn framework capitalizes on knowledge transfer to bolster its overall efficacy, streamlining the synthetic data generation process. In Section 4.3, we illustrate knowledge transfer in generative models using a pre-trained model based on adult male data (Kohavi et al., 1996), subsequently fine-tuned with adult female data for downstream analysis. As demonstrated in Figures 6, 7, and Table 3, the fine-tuned model adeptly captures the data distribution, even with a limited size of raw samples.
3 Syn-Slm: Streamlined Approach
The Syn Framework enables unsupervised synthetic data generation, mirroring raw data distributions. It is adept at tackling statistical challenges in both unsupervised and supervised realms. It can also derive $\tilde{F}$ directly from some generative models such as normalizing flows (Kingma and Dhariwal, 2018). Subsequently, we will explore a streamlined method, termed Syn-Slm, which bypasses synthetic data generation.
3.1 Synthetic Data Distribution
[…]
3.2 Statistical Methodologies
Supervised learning aims to predict an outcome based on a predictor vector $\bm{X}$. Consider a one-dimensional outcome variable $Y$ and define $\bm{Z}=(Y,\bm{X})$. In this scenario, generative models, such as diffusion models (Sohl-Dickstein et al., 2015; Ho, Jain and Abbeel, 2020) and the Reversible Generative Models (Kingma and Dhariwal, 2018), not only can generate synthetic data conditioning on $\bm{X}$ but also provide the conditional CDF of synthetic data $\tilde{F}_{Y|\bm{X}}$ as an estimate of the conditional cumulative distribution function (CDF) $F_{Y|\bm{X}}$. From this, one can deduce an optimal prediction function $f^{o}$ by minimizing the expected loss: $f^{o}=\inf_{f}\operatorname{E}L(Y,f(\bm{X}))$. Here, $L(\cdot,\cdot)$ symbolizes a loss function. For instance, if $l$ is the 00-$1$ loss in binary classification, then $f^{o}=F_{Y|\bm{X}}-1/2$, and its Syn-Slm estimate is given by $\tilde{f}^{o}=\tilde{F}_{Y|\bm{X}}-1/2$. An example is available in Section 4.1.
@@ 4.1 Sentiment Analysis
This subsection presents sentiment classification applied to the benchmark dataset, IMDB (Maas et al., 2011). This task involves assigning emotions expressed in the text into positive or negative sentiments based on the opinions reflected in each text. The dataset comprises $50,000$ polarized movie reviews, categorized as “positive” or “negative” sentiments. These labels correspond to movie scores below four or above seven out of ten, where no movie has more than 30 reviews to prevent significant class imbalance. Here, we use $49,000$ of these reviews as our training data, reserving $1,000$ reviews for testing. We compare Syn’s generative approach against its conventional counterpart in a downstream prediction task, utilizing three state-of-the-art models, GPT-3.5, DistilBERT, and LSTM models (Chen et al., 2023).
GPT-3.5 functions primarily as a text completion model, predicting the succeeding token as the sentiment label. Although essentially a completion model, GPT-3.5 is a conditional generative model that aligns with Syn, which we adapt for predictive tasks. We fine-tune GPT-3.5 with the text-embedding-Ada-002 configuration, adhering to OpenAI’s recommended procedures [^1] OpenAI GPT fine-tuning: https://platform.openai.com/docs/guides/fine-tuning.
In contrast, DistilBERT generates a fixed-size embedding of a review, which is then relayed to an appended classification head to deduce sentiment likelihood. Unlike GPT-3.5’s token generation approach, DistilBERT’s technique aligns more closely with Syn-Slm for supervised tasks. We fine-tune DistilBERT using the Distilbert-base-Uncased model from HuggingFace [^2] HuggingFace: https://huggingface.co/distilbert-base-uncased. Our tuning regimen includes a batch size of 16, a span of 10 epochs, and the Adam optimizer with standard decay parameters, setting the learning rate to $1\times 10^{-5}$.
Additionally, we train a traditional LSTM model from scratch, eschewing prior knowledge. Like DistilBERT, the LSTM processes an embedding and feeds it into a classification head, rendering it a predictive model. For our LSTM configuration, we modified code from a pertinent Kaggle notebook [^3] Kaggle: https://www.kaggle.com/code/pawan2905/imbd-sentiment-analysis-using-pytorch-lstm. We employ identical model structures and hyperparameters across our custom-split datasets.
Table 1 compares GPT-3.5, DistilBERT, and LSTM in seven performance metrics. The extensive collection of pre-trained models likely contributes to GPT-3.5’s outstanding performance. On the other hand, LSTM’s underperformance stems from its inability to transfer knowledge. Knowledge transfer plays a crucial role in model performance.
@@ 4.2 Prediction for Structured Data
This subsection investigates the generational phenomenon and challenges associated with enhancing the precision of gradient-boosting for regression and classification (Breiman, 1997; Friedman, 2002). It also focuses on the implications for the quality of synthetic data generation. Within the Syn framework, we designate the boosting method tailored for synthetic data as Syn-Boost. Despite the surge of diverse predictive models, the capabilities of the Syn framework in predictive modeling remain largely untapped. To highlight this potential, we draw contrasts between Syn-Boost and its traditional supervised counterparts: specifically, the boosting algorithm — CatBoost (Dorogush, Ershov and Gulin, 2018) and FNN — a fully connected neural network that leverages insights from a pre-trained model. Syn-Boost presents a strategy to harness knowledge transfer in boosting, effectively addressing the transfer learning challenge inherent to the boosting method.
@@ 4.2.1 Real-Benchmark Examples
To closely emulate real-world scenarios, we investigate situations where available pre-trained models have incorporated insights from relevant studies. For this study, we utilize five classification and three regression benchmark datasets (Kotelnikov et al., 2023), each encompassing three subsets: pre-training, raw, and test data. We train pre-trained models using the pre-training data, facilitating effective knowledge transfer. For Syn-Boost and FNN, we exclude the pre-training data from our raw set, which is then used for training and fine-tuning both methods. The test data, on the other hand, is reserved exclusively for performance evaluation. A detailed description of these datasets can be found in Table 4.
In the Syn-Boost framework, we utilize a tabular diffusion model, TDM (Kotelnikov et al., 2023), to generate synthetic data of mixed types that closely match the distribution of the original data. TDM employs multinomial and Gaussian diffusion processes to simulate categorical and continuous attributes. The procedure starts by training a TDM model [^4] We train TDMs using a single TITAN RTX GPU. on pre-existing data and then fine-tuning it with raw data. Subsequently, we use CatBoost on the synthetic data of size $m$, created by TDM for classification and regression tasks. To identify the best $m$ for Syn-Boost’s synthetic data, we evaluate the error relative to the synthetic-to-raw data ratio, ranging from $1$ to $30$, with a step size of $1$.
@@
Figure 2 underscores the significant contribution of Syn in bolstering CatBoost’s efficacy in classification and regression. While there is a minor boost in “FB comments”, the extent of improvement via Syn-Boost is diverse. Classification and regression enhancements span 0.6% to 17.4% and 0.03% to 12.3%, respectively, against CatBoost’s baseline rather than the Bayes error. Should the Bayes error have been accessible, we anticipate even higher percentage improvements. The magnitude of these enhancements varies by scenario. For instance, datasets like “Gesture” and “House”, which have a larger predictor count, show more significant leaps. In contrast, datasets like “Adult” and “California” with fewer predictors demonstrate modest gains.
@@
Figure 2: Comparative error analysis of Syn-Boost (blue), CatBoost (grey), and FNN (red) utilizing transfer learning on Eight Benchmarks (Kotelnikov et al., 2023) across different synthetic-to-raw ratios. The blue and red stars show the Syn-Boost sample size and the pre-training size for training FNN. Misclassification rates and RMSE measure classification and regression performance, respectively. A spline-smoothed curve with a 95% confidence interval of the fitted value (Hastie et al., 2009) is depicted in blue shades, highlighting error variation.
Syn’s robust performance primarily stems from the generative capabilities of diffusion models coupled with the application of knowledge transfer. These elements enhance the generative model’s generation accuracy by accurately estimating the distribution of raw data over low-dimensional manifolds (Oko, Akiyama and Suzuki, 2023). However, it is crucial to recognize that the Syn framework’s success also hinges on thoughtful modeling and predictor selection.
@@
In a supervised setting, Syn-Boost, which employs CatBoost on synthetic data, typically outperforms FNN, except in the “FB comments” and “Abalone” datasets. When the generation errors are modest, the performance edge of Syn-Boost over FNN spans from 11.1% to 14.6% in classification and 7.2% to 16.3% in regression. The reduced performance on the “FB comments” and “Abalone” datasets, with a decline ranging from 1.4% to 6.6%, is chiefly attributed to significant generation errors from the pre-trained TDM (Kotelnikov et al., 2023). In both scenarios, Syn-Boost may not surpass CatBoost when $m=n$. A similar phenomenon also occurs for Decision Tree, Random Forest, and Logistic Regression without knowledge transfer (Kotelnikov et al., 2023). We speculate that the generation error in the “FB comments” dataset arises from the model architecture’s inability to handle large pre-training instances, while the “Abalone” dataset’s underperformance could be due to insufficient pre-training size. Notably, while FNN outperforms CatBoost in the “Abalone”, “FB comments”, and “Gesture” datasets, it lags in others. These findings highlight Syn-Boost as a strong competitor against the well-established predictive model FNN.
@@
This case study highlights the effectiveness of the Syn framework in enhancing statistical accuracy through synthetic data generation. It also explains the results of Kotelnikov et al. (2023) regarding the potential pitfalls in a prediction task when employing synthetic data to train machine learning models, a concern mentioned in the Introduction. Low-fidelity synthetic data, resulting from substantial generation errors, can negatively impact statistical accuracy. The study suggests that employing knowledge transfer from relevant studies is a strategy to reduce generation errors.
@@ 4.2.2 Simulation
To investigate how generation errors impact the efficacy of Syn-Boost, we conducted simulations with access to ground truth data. We consider a model that closely mimics real benchmark examples:
@@
where $\bm{X}=(\bm{X}_{1},\ldots,\bm{X}_{7})$ is uniformly distributed over $[0,1]^{7}$ ($\text{Uniform}(0,1)^{7}$) and $\bm{\epsilon}$ follows a normal distribution with mean zero and standard deviation $0.2$, $N(0,.2^{2})$. In (4), we generate a dataset of 700 samples, dividing it into 500 for training and 200 for validation. To demonstrate the impact of effective versus ineffective generators on downstream tasks, we pre-train a tabular diffusion model (TDM, Kotelnikov et al. (2023)) with two sizes, $1000$ and $5000$. It is noteworthy that pre-trained models typically use considerably larger training sizes. To evaluate the distribution discrepancy between raw and synthetic samples, we employ the 2-Wasserstein distance, defined as their distributional distance and determined by solving an optimal transport problem using appropriate metrics [^5] https://pythonot.github.io/quickstart.html#computing-wasserstein-distance, as detailed in Table 2 for reference.
@@
4.3 Feature Relevance for Tabular Data
This subsection concerns the relevance of features for predicting the outcome of the response variable $Y$ by a machine learner using a candidate feature vector $\bm{X}$.
Define the subvector $\bm{X}_{S}$ by $\bm{X}_{S}=(X_{j}:j\in S)$, where $S$ is a subset of the features. Our objective is significance testing of $\bm{X}_{S}$ in its functional relevance to $Y$. To assess the influence of $\bm{X}_{S}$, we use the differenced risk $R(f^{*})-R(f_{S^{c}}^{*})$. Here, $f_{S^{c}}=f(\bm{X}_{S^{c}})$, and $f^{*}$ and $f^{*}_{S^{c}}$ are the optimal prediction functions in the population, defined as $f^{*}=\operatorname*{\arg\min}_{f}R(f)$ and $f^{*}_{S^{c}}=\operatorname*{\arg\min}_{f_{S^{c}}}R(f_{S^{c}})$. The risks, $R(f)$ and $R(f_{S^{c}})$, are given by $R(f)=\operatorname{E}\big{(}L(f(\bm{X}),Y)\big{)}$ and $R(f_{S^{c}})=\operatorname{E}\big{(}L(f(\bm{X}_{S^{c}}),Y)\big{)}$, where $\operatorname{E}$ represents the expectation over randomness. Now, we introduce the null and its alternative hypotheses $H_{0}$ and $H_{a}$:
For Syn-Test, we adhere to Steps 1-4, as delineated in Section 2.3, to examine the relevance of feature set $\bm{X}_{S}$ to outcome $Y$, employing CatBoost (Dorogush, Ershov and Gulin, 2018) as the learning algorithm. Here, we engage a diffusion model, TDM (Kotelnikov et al., 2023), to engender synthetic data. Initially, we adapt the original test statistic from (Dai, Shen and Pan, 2022) to suit synthetic data as follows:
Figure 3: RMSE of Syn-Boost versus synthetic to raw data ratio in simulations (4). It compares small (blue, training size=1K) and large (red, training size=5K) generation errors and generation with (solid line) and without (dotted line) fine-tuning on raw data. The grey dashed and black solid lines represent the RMSE of CatBoost and the square root of Bayes error, respectively.

### 10 Mar 2024 v3

@@ 2.5 Generative Model and Knowledge Transfer
The Syn framework capitalizes on knowledge transfer to bolster its overall efficacy, streamlining the synthetic data generation process. For a visual representation of how a pre-trained generative model is fine-tuned on a specific dataset to facilitate knowledge transfer, refer to Figure 1. In Section 3.3, we illustrate knowledge transfer in generative models using a pre-trained model based on adult male data (Kohavi et al., 1996), subsequently fine-tuned with adult female data for downstream analysis. As demonstrated in Figures 3, 4, and Table 3, the fine-tuned model adeptly captures the data distribution, even with a limited size of raw samples.
@@ 3.1 Sentiment Analysis
This subsection presents sentiment classification applied to the benchmark dataset, IMDB (Maas et al., 2011). This task involves assigning emotions expressed in the text into positive or negative sentiments based on the opinions reflected in each text. The dataset comprises $50,000$ polarized movie reviews, categorized as “positive” or “negative” sentiments. These labels correspond to movie scores below four or above seven out of ten, where no movie has more than 30 reviews to prevent significant class imbalance. Here, we use $49,000$ of these reviews as our training data, reserving $1,000$ reviews for testing. We compare Syn’s generative approach against its conventional counterpart in a downstream prediction task, utilizing three models, GPT-3.5, DistilBERT, and LSTM models (Chen et al., 2023).
GPT-3.5 functions primarily as a text completion model, predicting the succeeding token as the sentiment label. Although essentially a completion model, GPT-3.5 is a conditional generative model that aligns with Syn-Slm, which we adapt for predictive tasks. In contrast, DistilBERT generates a fixed-size embedding of a review, which is then passed to an appended classification head to deduce sentiment likelihood. Unlike GPT-3.5’s token generation approach, DistilBERT’s technique aligns more closely with traditional predictive modeling with transfer learning for supervised tasks. Additionally, we train a traditional LSTM model from scratch, eschewing prior knowledge. Like DistilBERT, the LSTM processes an embedding and feeds it into a classification head, rendering it a predictive model.
Table 2 compares GPT-3.5, DistilBERT, and LSTM in six performance metrics. The extensive collection of pre-trained models likely contributes to GPT-3.5’s outstanding performance. On the other hand, LSTM’s underperformance stems from its inability to transfer knowledge. Knowledge transfer plays a crucial role in model performance. For details on training configurations, refer to the Supplementary Materials.
@@ 3.2 Prediction for Structured Data
This subsection investigates the generational phenomenon and challenges associated with enhancing the precision of gradient-boosting for regression and classification (Breiman, 1997; Friedman, 2002). It also focuses on the implications for the quality of synthetic data generation. Within the Syn framework, we designate the boosting method tailored for synthetic data as Syn-Boost. Despite the surge of diverse predictive models, the capabilities of the Syn framework in predictive modeling remain largely untapped. To highlight this potential, we draw contrasts between Syn-Boost and its traditional supervised counterparts: specifically, the boosting algorithm — CatBoost (Dorogush, Ershov and Gulin, 2018) and FNN — a fully connected neural network that leverages insights from a pre-trained model, both of which are traditional. Syn-Boost presents a strategy to harness knowledge transfer in boosting, effectively addressing the transfer learning challenge inherent to the boosting method.
@@ 3.2.1 Real-Benchmark Examples
To closely emulate real-world scenarios, we investigate situations where available pre-trained models have incorporated insights from relevant studies. For this study, we utilize five classification and three regression benchmark datasets (Kotelnikov et al., 2023), each encompassing three subsets: pre-training, raw, and test data. A detailed description of these datasets can be found in Table 6.
In the Syn-Boost framework, we utilize a tabular diffusion model, TDM (Kotelnikov et al., 2023), to generate synthetic data of mixed types that closely match the distribution of the original data. TDM employs multinomial and Gaussian diffusion processes to simulate categorical and continuous attributes. The procedure starts by training a TDM model on pre-existing data and then fine-tuning it with raw data. Subsequently, we use CatBoost on the synthetic data of size $m$, created by TDM for classification and regression tasks. To identify the best $m$ for Syn-Boost’s synthetic data, we evaluate the error relative to the synthetic-to-raw data ratio, ranging from $1$ to $30$, with a step size of $1$.
@@
Knowledge Transfer from Identical Distributions. For each dataset mentioned, we utilize pre-training data to train pre-trained models, enabling the downstream knowledge transfer on raw data with either Syn-Boost or FNN. We use the test data solely for evaluating performance. It is important to note that in this study, the pre-training and raw data used for fine-tuning share the same distribution.
Figure 2: Comparative error analysis of CatBoost, Syn-Boost, and FNN, with Syn-Boost and FNN applying transfer learning with the same distributions across eight benchmarks (Kotelnikov et al., 2023), measured at various synthetic-to-raw data ratios. The stars indicate the size of the pre-training data used to obtain pre-trained models and the tuned sample size for Syn-Boost. The performance for classification and regression tasks is measured by misclassification rate and RMSE, respectively. Point-wise standard errors, derived from smoothing spline models (Hastie et al., 2009), are also depicted to illustrate the variation in error.
Figure 2 underscores the contribution of Syn in bolstering CatBoost’s efficacy in classification and regression. While there is a minor boost in “FB comments”, the extent of improvement via Syn-Boost is diverse. Classification and regression enhancements span 0.6% to 17.4% and 0.03% to 12.3%, respectively, against CatBoost’s baseline. The magnitude of these enhancements varies by scenario. For instance, datasets like “Gesture” and “House”, which have a larger predictor count, show more significant leaps. In contrast, datasets like “Adult” and “California” with fewer predictors demonstrate modest gains.
@@
In a supervised setting, Syn-Boost, which employs CatBoost on synthetic data, typically outperforms FNN, except in the “FB comments” and “Abalone” datasets. When the generation errors are modest, the performance of Syn-Boost over FNN spans from 11.1% to 14.6% in classification and 7.2% to 16.3% in regression. The reduced performance on the “FB comments” and “Abalone” datasets, with a decline ranging from 1.4% to 6.6%, is chiefly attributed to non-negligible generation errors from the pre-trained TDM (Kotelnikov et al., 2023). In both scenarios, Syn-Boost may not surpass CatBoost when $m=n$. A similar phenomenon also occurs for other models including decision trees, random forests, and logistic regression without knowledge transfer (Kotelnikov et al., 2023). We speculate that the generation error in the “FB comments” dataset arises from the model architecture’s inability to handle large pre-training instances, while the “Abalone” dataset’s underperformance could be due to insufficient pre-training size. Notably, while FNN outperforms CatBoost in the “Abalone”, “FB comments”, and “Gesture” datasets, it lags in others. These findings highlight Syn-Boost as a strong competitor against well-established predictive models.
Knowledge Transfer Across Distinct Distributions. The Adult dataset (Kohavi et al., 1996), derived from the $1994$ census, includes data on $32,650$ males and $16,192$ females, featuring six numerical and eight nominal attributes such as age, work class, education years, marital status, weekly working hours, and native country. Our goal is to predict whether an adult female’s income surpasses \$50K annually using a generative model pre-trained on male data to facilitate knowledge transfer, despite their differing distributions but capitalizing on their similarities.
As Figure 3 depicts, the pronounced differences across genders exist in the distributions of categories like income, age, marital status, occupation, and relationship. A pertinent question arises: can adult male data augment the synthetic data generation and subsequent classification task on adult female data? To harness this knowledge transfer, we pre-train a TDM solely on the adult-male data with $32,650$ observations. We utilize a subset of size $1,350$ from the adult-female data as our raw data, and another independent subset of the same size as the test data for evaluation purposes.
Table 3: Distribution distance comparisons using FID-scores, 1- and 2-Wasserstein distances between the actual female sample and various synthetic samples. The terms ”raw,” ”pre-trained,” and ”fine-tuned” denote the raw data, synthetic data from a pre-trained model, and synthetic data from a fine-tuned pre-trained model, respectively. The evaluation sample size for ”Female (raw)” is $1,350$, while for ”Male (raw),” ”Male (pre-trained),” ”Female (pre-trained),” and ”Female (fine-tuned),” is $32,650$.
FID (Gaussian) 1-Wasserstein 2-Wasserstein
Female (raw) vs Male (raw) 1.971 1.968 2.125
Female (raw) vs Male (pre-trained) 2.051 1.967 2.127
Female (raw) vs Female (pre-trained) 0.457 1.292 1.536
Female (raw) vs Female (fine-tuned) 0.249 1.170 1.399
Figures 3 and 4 demonstrate that the synthetic female data produced by TDM aligns more with the adult-female dataset than with the adult-male dataset. This alignment is evident by the Fréchet Inception Distance (FID), which measures the distributional differences between the generated and raw data vectors under the Gaussian assumption. The 1- and 2-Wasserstein distances between the empirical distributions of the two datasets provide further evidence (note that FID is 2-Wasserstein distance under Gaussian assumption). Notably, as shown in Table 3, the male-focused pre-trained TDM, once fine-tuned with a smaller female dataset, crafts synthetic female data with a diminished margin of error compared to models trained solely on the adult-male data and adult-female data. This result confirms that leveraging pre-trained adult-male data with a somewhat distinct distribution can enhance the TDM’s generation precision for females. This empirical validation emphasizes the imperative of refining a pre-trained model to maximize knowledge transfer and achieve unparalleled accuracy.
Figure 3: Marginal distributions of datasets categorized as female, synthetic female, and male are illustrated, with legends arranged from top to bottom. Normalized bar and kernel density plots represent categorical and numerical features, respectively.
The result of Syn-Boost prediction is illustrated in Figure 5. We observe that simply increasing the synthetic size does not necessarily improve the prediction performance, as depicted at the beginning of the Syn-Boost tuning. The optimal empirical ratio is obtained at $m/n=23$, which is about the same size as the pre-training adult-male data. Moreover, Syn-Boost achieves $5.32\%$ improvement compared to CatBoost in misclassification error, and $11.00\%$ improvement compared to the FNN approach, indicating the potential of this method.
@@
Both case studies highlight the effectiveness of the Syn framework in enhancing statistical accuracy through synthetic data generation. Syn’s robust performance primarily stems from the generative capabilities of diffusion models coupled with the application of knowledge transfer. These elements enhance the generative model’s generation accuracy by accurately estimating the distribution of raw data over low-dimensional manifolds (Oko, Akiyama and Suzuki, 2023). However, it is crucial to recognize that the Syn framework’s success also hinges on thoughtful modeling and predictor selection. The case studies also explain the results of Kotelnikov et al. (2023) regarding the potential pitfalls in a prediction task when employing synthetic data to train machine learning models, a concern mentioned in the Introduction. Low-fidelity synthetic data, resulting from substantial generation errors, can negatively impact statistical accuracy. The study suggests that employing knowledge transfer from relevant studies is a strategy to reduce generation errors.
Figure 5: Comparative error analysis of CatBoost, Syn-Boost, and FNN, with Syn-Boost and FNN utilizing transfer learning with distinct distributions on the Adult-Female dataset (Kohavi et al., 1996), with Adult-Male data serving as pre-training data, across various synthetic-to-raw data ratios. Stars indicate the pre-training data size and the tuned sample size for Syn-Boost. The vertical bars, calculated using smoothing spline models (Hastie et al., 2009), represent the pointwise standard error.
@@ 3.2.2 Simulation
To investigate how generation errors impact the efficacy of Syn-Boost, we conducted simulations with access to ground truth data. We consider a regression model:
@@
where $\bm{X}=(\bm{X}_{1},\ldots,\bm{X}_{7})$ is uniformly distributed over $[0,1]^{7}$ ($\text{Uniform}(0,1)^{7}$) and $\bm{\epsilon}$ follows a normal distribution with mean zero and standard deviation $0.2$, $N(0,.2^{2})$. In (4), we generate a dataset of 700 samples, dividing it into 500 for training and 200 for validation. To demonstrate the impact of effective versus ineffective generators on downstream tasks, we pre-train tabular diffusion models (Kotelnikov et al., 2023)) with two sizes, $1000$ and $5000$. It is noteworthy that pre-trained models typically use considerably larger training sizes. To evaluate the distribution discrepancy between raw and synthetic samples, we employ the 2-Wasserstein distance, defined as their distributional distance and determined by solving an optimal transport problem using appropriate metrics [^2] https://pythonot.github.io/quickstart.html#computing-wasserstein-distance, as detailed in Table 4 for reference.
@@
Figure 6: RMSEs of Syn-Boost as a function of the synthetic to raw data ratio in simulations described in (4). It contrasts scenarios with small (pre-training size of 1,000) and large (pre-training size of 5,000) generation errors, as well as the generation with (dotted lines) and without (solid lines) fine-tuning on raw data. The grey dashed lines denote the RMSE of CatBoost, while the black solid lines indicate the square root of the Bayes error.
3.3 Feature Relevance for Tabular Data
This subsection concerns testing the relevance of features for predicting the outcome of the response variable $Y$ by a machine learner using a candidate feature vector $\bm{X}$.
Define the subvector $\bm{X}_{S}$ by $\bm{X}_{S}=(X_{j}:j\in S)$, where $S$ is a subset of the features. Our objective is significance testing of $\bm{X}_{S}$ in its functional relevance to $Y$. To assess the influence of $\bm{X}_{S}$, we use the differenced risk $R(f^{*})-R(f_{S^{c}}^{*})$. Here, $f_{S^{c}}=f(\bm{X}_{S^{c}})$, and $f^{*}$ and $f^{*}_{S^{c}}$ are the optimal prediction functions in the population, defined as $f^{*}=\operatorname*{\arg\min}_{f}R(f)$ and $f^{*}_{S^{c}}=\operatorname*{\arg\min}_{f_{S^{c}}}R(f_{S^{c}})$. The risks, $R(f)$ and $R(f_{S^{c}})$, are given by $R(f)=\operatorname{E}\big{(}L(f(\bm{X}),Y)\big{)}$ and $R(f_{S^{c}})=\operatorname{E}\big{(}L(f(\bm{X}_{S^{c}}),Y)\big{)}$, where $\operatorname{E}$ represents the expectation. Now, we introduce the null and its alternative hypotheses $H_{0}$ and $H_{a}$:
For Syn-Test, we adhere to Steps 1-4, as delineated in Section 2.3, to examine the relevance of feature set $\bm{X}_{S}$ to outcome $Y$, employing CatBoost (Dorogush, Ershov and Gulin, 2018) as the learning algorithm. Here, we engage a diffusion model, TDM (Kotelnikov et al., 2023), to generate synthetic data. Initially, we adapt the original test statistic from (Dai, Shen and Pan, 2024) to suit synthetic data as follows: